### v2

@@ Abstract
In this article, we discuss the ubiquity of Big Data and machine learning in society and propose that it evinces the need of further investigation of their fundamental limitations. We extend the “too much information tends to behave like very little information” phenomenon to formal knowledge about lawlike universes and arbitrary collections of computably generated datasets. This gives rise to the simplicity bubble problem, which refers to a learning algorithm equipped with a formal theory that can be deceived by a dataset to find a locally optimal model which it deems to be the global one. In the context of lawlike (computable) universes and formal learning systems, we show that there is a ceiling above which formal knowledge cannot further decrease the probability of zemblanitous findings, should the randomly generated data made available to the formal learning system be sufficiently large in comparison to their joint complexity. We also argue that this is an epistemological limitation that may generate unpredictable problems in digitally intermediated societies.
@@ 1 Introduction
The large amount of collected data coupled with the success of AI algorithms is rapidly increasing the pervasiveness of computational intermediation of humans and institutional relationships. Some enthusiasts of the use of new technologies for data collection and data analysis argue that the access to ever-increasing volumes of diverse data has made the formulation of explanatory hypotheses and the development of scientific theories, along with the elaboration of causal models, obsolete Anderson (2008). In this sense, Big Data enthusiasts imply and believe that the data analysis is capable of making serendipitous discoveries Cavassane (2022): findingsthat are unpredictable, random, or appear to be made out of the blue with respect to the data analyst’s scientific skills or current state of knowledge, but nevertheless happen to be fortunate in the end. Being fortuitous, serendipity would enable one to solve a previously stated problem in an unforeseen manner or develop irreducibly new theoretical findings Cavassane (2022).
However, in the context of Big Data analysis, we argue that a zemblanatious finding should be more prominent than a serendipitous one. Zemblanity is the polar opposite of serendipity Boyd (1998). While serendipity refers to an unexpected or unpredictable event that presents an opportunity for a fortunate discovery, zemblanity refers to a result that should have been expected/predictable which reveals an underlying problem or negative consequence in a given model or theory. It is not a discovery at all, but rather a systemic flaw or negative aspect that the observer is either intentionally or unintentionally overlooking, not acknowledging in the results’ statements, or somehow refusing to understand. This is because both in principle and in practice a zemblanitous phenomenon would be predictable by the data analyst, being one of the possible direct implications of (or one that can be inferred from) the current theory and state of knowledge. Unfortunately, these deleterious or problematic properties are often only revealed when these conclusions are put into practice in society and cause societal conflict and damage.
@@
On the one hand, serendipitous findings occur by chance, or are unpredictable, but are valuable with respect to solving a previously stated problem by introducing new and promising hypotheses or lines of investigation that can be used to develop a new theory or extend previous theories Merton (1948). On the other hand, a zemblanitous finding is a result of biases or errors that will necessarily appear in the model. Thus, since the term “zemblanity” means the complementary antonym of “serendipity”, a zemblanitous finding is not the product of chance, but actually inherent to these Big Data technologies. [^1] A second type of zemblanity could consist not in the complementary antonym of “serendipity”, but in its gradative antonym, that is, meaning not a necessary result, but one which, though expected to eventually happen, is also triggered by a random event. We could thus admit intermediary degrees between both notions, with an unexpected and valuable finding generated by “good luck” on one side of the scale, and an expected and negative result generated by “bad luck” on the other side of the scale. We will deal with that alternative definition in greater detail in another work, however.
@@
The phenomenon of “too much information tends to behave like very little information” is in consonance with on aspect of machine learning that developers of the technology have grappled with since the inception of AI: avoiding overfitting and guaranteeing that the learner will not get stuck in local optima. In this article, we aim to show that these issues are the result of a systemic problem: that there is a deceiving phenomenon inherent to these systems that result from a fundamental limit of formal knowledge systems.
@@
Overfitting refers to the case in which the error on the dataset against which the trained model candidate is tested is considered to be relatively higher than the training error, which was employed (and previously chosen) as the error bounds or margins in the training phase so the learner would classify a model candidate as successful with regards to the desired task. On the other hand, underfitting occurs when the value of the loss function (i.e., the function that calculates how much a model candidate diverges from the optimal one, or from the target values) retrieved from the training dataset is considered to be higher than that training error. Underfitting can be easily solved by increasing the number, or the complexity, of the parameters, hyperparameters, layers of nodes in a neural network, etc. Overfitting demands much more sophisticated methods along with human supervision and agency. These methods must effect a procedure, such as regularization, that adapts or translates some heuristics, inspired by Occam’s razor or a bias toward simplicity, into a successful and tractable computational implementation of an algorithm that performs the desired task within the required error bounds Goodfellow et al. (2016); Witten et al. (2017). By embedding such a bias toward simplicity into the regularization phase itself, recent results have shown that models that yield a perfect fit with a minimum training error for every data point in the training set tend to naturally have greater algorithmic complexity so that overfitting can be avoided by design Hernández-Orozco et al. (2021).
@@
The no free lunch theorem Wolpert and Macready (1997) says that no learning algorithm is any better than any other equally likely learning algorithm Goodfellow et al. (2016). When averaging out all possible optimization problems, this means that for every learner there will exist a task at which the learner succeeds as much as another learner that already fails at it. This implies a theoretical limitation where one cannot guarantee that there is a general method in which the learner avoids getting stuck in locally optimal models, i.e., that the learner never approximates the globally optimal solution within the desired error margins. In practice, one can assume distributions in real-world data gathering that favor simpler models, thus avoiding the condition of performing the uniform average. This bias toward simplicity constitutes a way out of the conditions for the no free lunch theorem Goodfellow et al. (2016); Witten et al. (2017); Scholkopf et al. (2021) because in practice one can reduce the scope of the problem by imposing a priori constraints on the possible model candidates.
@@
Such limitations are systemic to any formal-theoretic or computational method that learns through experience in arbitrarily large lawlike universes. By experience [^2] Note that in the “metaphenomenological” model proposed in Winter (2020), the concept of the digital experience is originally defined in reverse: it is an algorithm that takes as input a subject/model and transforms it. This theoretically allows for arbitrarily large mutations that could overcome the problem of getting trapped in local optima, but only with a limited use of an oracle to avoid non-halting results similar to Chaitin’s metabiological modelChaitin (2012). We derived this simplified definition as it more accurately reflects the way machine learning algorithms currently function., we mean a set of events input to a learning algorithm equipped with a formal theory Winter (2020) A lawlike universe follows a nomic assumption about its own structure: that it can be explained or generated by a finite set of generative rules that are much simpler than the whole set of events or outcomes. In particular, this is the case of any possible universe that is in principle computable Zenil (2013); Abrahão and Zenil (2022)—which holds even if one might never know the universe’s original generative algorithm.
@@ 2 The simplicity bubble problem
As a result, large enough datasets can deceive learning algorithms into accurately finding a (locally) optimal solution (which the learning algorithm may return as being a global optimum candidate) while never being able to actually find a globally optimal solution for the whole universe of possible events. This is a simplicity bubble effect (SBE) Abrahão et al. (2023b): the phenomenon where the underlying data generating process (mechanism or external source) is trapping an algorithm (or a formal theory) into a simplicity bubble because the data made available by the generating process is complex enough to ensure that any model proposed by the algorithm (or formal theory) is insufficiently complex to successfully approximate a globally optimal model that actually corresponds to any fresh data that the algorithm may receive. The SBE has to do with a mathematical phenomenon that appears from a (algorithmic) complexity-centered interplay between data and algorithms. Another form of interpreting this idea is that the deception originates from a wide enough gap between the algorithmic information of the underlying generative model, laws, or principles of the external source of the events or phenomena and the algorithmic information of the formal learning systems, i.e., the observer. This gap is guaranteed when the available dataset is sufficiently large. As more and more data becomes available, its underlying generating process (or mechanism) likely becomes more complex than the observer’s tools or formal knowledge themselves Abrahão et al. (2023b). Only by adding a larger amount of (unpredictable) fresh data, i.e., new phenomena, not yet available to the observer in first place would the observer hypothetically be capable of “bursting” this bubble, which in turn would make the observer “realize” that the optimum it did proposed to be the global one actually is not the global optimum.
@@ 2.1 No formal learning system can rule out demons in justified theories for lawlike universes
Let an hypothetical abstract entity, i.e., a ‘demon’, in a thought experiment correspond to a sufficiently large dataset (or to the underlying generative and governing agent for the set of every possible phenomena) which is an unpredictable deceiver that triggers the SBE for a given observer, i.e., a formal learning system. The above epistemological limitation is essentially a lower bound for the probability (given any arbitrary formal theory that defines the arbitrarily chosen probability measure) that such a demon exists. As the set of known and explainable phenomena becomes arbitrarily large, if Conditions I and II are satisfied, no one can rule out the possibility of those demons by endlessly decreasing the probability of their existence.
@@ 3 Interplay between Big Data, formal learning systems, and humanity
Leonelli proposes the notion of data journeys Leonelli (2014, 2023) to account for the material aspects, local idiosyncrasies, experimental settings, inferential processes, among other elements which influences the results in data-intensive research related to the processes of data achievement, integration, and use. Data journeys involve the process of de-contextualization of data from their local origin, such that it becomes digital and integrated in structured or semi-structured datasets that can be retrieved in other domains, allowing the possibilities of data reuse, reanalysis, recombination, repurposing, and repositioning Leonelli (2014); Collmann et al. (2016).
@@
An example of a data journey is the process of retrieval of information about amino acid sequences stored in genomic data infrastructures, in order to design new forms of protein structures using freely available software for academic research Anishchenko et al. (2021). The storage of protein data in massive datasets, such as Uniprot1, requires the process of standardization of the uploaded forms of proteins, being experimentally achieved or computationally designed. In addition, the dataset can impose some limitation in the retrieval of information. For example, a maximum of one hundred sequences per search. In the endeavor to cover as many species as possible, there could also be biases which favor the species that are well studied over fewer known organisms. Other kinds of biases may arise in the analysis phase itself as a limitation of statistics-based methods as is demonstrated in Uthamacumaran et al. (2023) when one tries to distinguish living from non-living matter.
@@
As presented in Cavassane et al. (2023), a potential problem illuminated by data journeys is the growing reliance on acknowledgeable datasets for decision-making processes in a software model which might be used to inform further research whose results might, in a feedback loop, serve as input to the data infrastructure. Our idea here is not to disregard the epistemic value of data integration, data modelling, and data retrieval, but to shed light on possible challenges of the growing automation of scientific practices by reflecting upon decision-making processes involved in data-journeys. These challenges and steps toward solving them are discussed in Zenil et al. (2023b), in which new theoretical and technological tools are proposed in order to deal with the dynamics between AI-driven scientific discoveries and human scientists themselves once both will belong to the same knowledge production system in a closed-loop manner.
@@
In summary, data journeys constitute dynamic structures that, in association with computational tools, might guide scientific research and decision-making processes, promising new forms of scientific discovery. Once suitable for data journeys, data can grow in meaning by the process of recontextualization, with their original purpose being revised to fit a different research question or to redirect the course of investigation. Similar to the framework given by algorithmic information dynamics (AID) Zenil et al. (2020); Abrahão and Zenil (2022); Zenil et al. (2019a, 2018, b) to study causal inference without the a priori separation between observers and objects, data journeys contrast the so-called ‘objective-oriented’ view of traditional scientific practice. One in which, by avoiding human interference and potential mistakes in scientific practice, we would eventually achieve neutral scientific results. This argument is succinctly summarized by Leonelli: “the paradox consists of the observation that, despite their epistemic value as ‘given’, data are clearly made” Leonelli (2014). Instead of this paradox suggesting an inexorable limit that science should acknowledge, novel methodologies should embed this paradigm shift in order to allow the investigation of both the objects and the observers themselves as interacting systems from the beginning and as part of the same theoretical and empirical framework Zenil et al. (2020); Abrahão and Zenil (2022); Zenil et al. (2023b).
@@
If we disregard, for the sake of the argument, the potential difficulties described above and imagine that a given dataset provides a reliable source for scientific evidence, the question to be reflected upon would be about the possibility to mechanically discriminate patterns which might unveil causal relationships among variables of interest from spurious correlations Floridi ; Smith (2020); Kitchin (2014); Calude and Longo (2017). A key criterion for the detection of causal relationships among variables is to discover the directionality of these variables, i.e., to uncover a causal chain (i.e., that rain causes the soil to be wet and not vice versa), or to show that two variables are the effect of the same cause (for instance, that thunder and lightning are caused by electric discharge). This reflects the Reichenbach principle: if two properties or kinds of events are probabilistically dependent, then they are causally connected in the sense that either one is a cause of the other (or vice versa), or both are effects of a common cause (where X is a cause of Y iff there leads a path of causal arrows from X to Y) Schurz (2008); Glymour et al. (1991).

### v3

@@ Abstract
In this article, we discuss the ubiquity of Big Data and machine learning in society and propose that it evinces the need of further investigation of their fundamental limitations. We extend the “too much information tends to behave like very little information” phenomenon to formal knowledge about lawlike universes and arbitrary collections of computably generated datasets. This gives rise to the simplicity bubble problem, which refers to a learning algorithm equipped with a formal theory that can be deceived by a dataset to find a locally optimal model which it deems to be the global one. In the context of lawlike (computable) universes and formal learning systems, we show that there is a ceiling above which formal knowledge cannot further decrease the probability of zemblanitous findings, should the randomly generated data made available to the formal learning system be sufficiently large in comparison to their joint complexity. Zemblanity, the opposite of serendipity, is defined by an undesirable but expected finding that reveals an underlying problem or negative consequence in a given model or theory, which is in principle predictable in case the formal theory contains sufficient information. We also argue that this is an epistemological limitation that may generate unpredictable problems in digitally intermediated societies.
@@ 1 Introduction
The large amount of collected data coupled with the success of AI algorithms is rapidly increasing the pervasiveness of computational intermediation of humans and institutional relationships. Some enthusiasts of the use of new technologies for data collection and data analysis argue that the access to ever-increasing volumes of diverse data has made the formulation of explanatory hypotheses and the development of scientific theories, along with the elaboration of causal models, obsolete Anderson (2008). In this sense, Big Data enthusiasts imply and believe that the data analysis is capable of making serendipitous discoveries Cavassane (2022): findings that are unpredictable, random, or appear to be made out of the blue with respect to the data analyst’s scientific skills or current state of knowledge, but nevertheless happen to be fortunate in the end. Serendipity combines both being an accident and the recognition of such accident as a valuable source for knowledge acquisition Merton (2000). Being fortuitous, serendipity would enable one to solve a previously stated problem in an unforeseen manner or develop irreducibly new theoretical findings Cavassane (2022).
However, in the context of Big Data analysis, we argue that a zemblanatious finding should be more prominent than a serendipitous one. Zemblanity is the polar opposite of serendipity Boyd (1998). While serendipity refers to an unexpected or unpredictable event that presents an opportunity for a fortunate discovery, zemblanity refers to a result that should have been expected/predictable which reveals an underlying problem or negative consequence in a given model or theory. It is not a discovery at all, but rather a systemic flaw or negative aspect that the observer is either intentionally or unintentionally overlooking, not acknowledging in the results’ statements, or somehow refusing to understand. This is because, both in principle and in practice, a zemblanitous phenomenon would be predictable by the data analyst, being one of the possible direct implications of (or one that can be inferred from) the current theory and state of knowledge. Unfortunately, these deleterious or problematic properties are often only revealed when these conclusions are put into practice in society and cause societal conflict and damage.
@@
On the one hand, serendipitous findings occur by chance, or are unpredictable, but are valuable with respect to solving a previously stated problem by introducing new and promising hypotheses or lines of investigation that can be used to develop a new theory or extend previous theories Merton (1948). On the other hand, a zemblanitous finding is a result of biases or errors that will necessarily appear in the model. Thus, since the term “zemblanity” means the complementary antonym of “serendipity”, a zemblanitous finding is not the product of chance, but actually inherent to these Big Data technologies. [^1] A second type of zemblanity could consist not in the complementary antonym of “serendipity”, but in its gradative antonym, that is, meaning not a necessary result, but one which, though expected to eventually happen, is also triggered by an unexpected event. One could thus admit intermediary degrees between both notions, with an unexpected and valuable finding generated by an accident with potentially positive outcomes (since serendipity can happen by accident and sagacity Walpole ; Andel (1994), or at the intersection of chance and wisdom Copeland (2019)) on one side of the scale; and an expected and negative result generated by an accident with potentially negative outcomes on the other side of the scale. Further investigation on this alternative definition is a necessary future work.
@@
The phenomenon of “too much information tends to behave like very little information” is in consonance with on aspect of machine learning that developers of the technology have grappled with since the inception of AI: avoiding overfitting and guaranteeing that the learner will not get stuck in local optima. In this article, we aim to show that these issues are the result of a systemic problem: that there is a deceiving phenomenon inherent to these systems that results from a fundamental limit of formal knowledge systems.
@@
Overfitting refers to the case in which the error on the dataset against which the trained model candidate is tested is considered to be relatively higher than the training error, which was employed (and previously chosen) as the error bounds or margins in the training phase so the learner would classify a model candidate as successful with regards to the desired task. On the other hand, underfitting occurs when the value of the loss function (i.e., the function that calculates how much a model candidate diverges from the optimal one, or from the target values) retrieved from the training dataset is considered to be higher than that training error. Underfitting can be easily solved by increasing the number, or the complexity, of the parameters, hyperparameters, layers of nodes in a neural network, etc. Overfitting demands much more sophisticated methods along with human supervision and agency. These methods must effect a procedure, such as regularization, that adapts or translates some heuristics, inspired by Occam’s razor or a bias toward simplicity, into a successful and tractable computational implementation of an algorithm that performs the desired task within the required error bounds Goodfellow et al. (2016); Witten et al. (2017). By embedding such a bias toward simplicity into the regularization phase itself, recent results have shown that models yielding a perfect fit with a minimum training error for every data point in the training set tend to naturally have greater algorithmic complexity so that overfitting can be avoided by design Hernández-Orozco et al. (2021).
@@
The no free lunch theorem Wolpert and Macready (1997) says that no learning algorithm is any better than any other equally likely learning algorithm Goodfellow et al. (2016). When averaging out all possible optimization problems, this means that for every learner there will exist a task at which the learner succeeds as much as another learner that already fails at it. This implies a theoretical limitation where one cannot guarantee that there is a general method in which the learner avoids getting stuck in locally optimal models. That is, one cannot guarantee that there is a general method in which the learner can always approximate the globally optimal solution within the desired error margins. In practice, one can assume distributions in real-world data gathering that favor simpler models, thus avoiding the condition of performing the uniform average. This bias toward simplicity constitutes a way out of the conditions for the no free lunch theorem Goodfellow et al. (2016); Witten et al. (2017); Scholkopf et al. (2021) because in practice one can reduce the scope of the problem by imposing a priori constraints on the possible model candidates.
@@
Such limitations are systemic to any formal-theoretic or computational method that learns through experience in arbitrarily large lawlike universes. By experience [^2] Note that in the “metaphenomenological” model proposed in Winter (2020), the concept of the digital experience is originally defined in reverse: it is an algorithm that takes as input a subject (i.e., in our present case, a model) and always transforms it into something else. This theoretically would allow the subject to overcome the problem of getting trapped in local optima, but only with a limited use of an oracle to avoid non-halting results similar to Chaitin’s metabiological model Chaitin (2012). The definition employed in this article so as to more accurately reflect the way machine learning algorithms currently function., we mean a set of events input to a learning algorithm equipped with a formal theory Winter (2020). A lawlike universe follows a nomic assumption about its own structure: that it can be explained or generated by a finite set of generative rules that are much simpler than the whole set of events or outcomes Abrahão and Arroyo (2023). In particular, this is the case of any possible universe that is in principle computable Zenil (2013); Abrahão and Zenil (2022)—which holds even if one might never know the universe’s original generative algorithm.
@@ 2 The simplicity bubble problem
As a result, large enough datasets can deceive learning algorithms into accurately finding a (locally) optimal solution (which the learning algorithm may return as being a global optimum candidate) while never being able to actually find a globally optimal solution for the whole universe of possible events. This is a simplicity bubble effect (SBE) Abrahão et al. (2023b): the phenomenon where the underlying data generating process (mechanism or external source) is trapping an algorithm (or a formal theory) into a simplicity bubble because the data made available by the generating process is complex enough to ensure that any model proposed by the algorithm (or formal theory) is insufficiently complex to successfully approximate a globally optimal model that actually corresponds to any fresh data that the algorithm may receive. The SBE has to do with a mathematical phenomenon that appears from a (algorithmic) complexity-centered interplay between data and algorithms. Another form of interpreting this idea is that the deception originates from a wide enough gap between the algorithmic information of the underlying generative model, laws, or principles of the external source of the events or phenomena and the algorithmic information of the formal learning systems, i.e., the observer. This gap is guaranteed when the available dataset is sufficiently large. As more and more data becomes available, its underlying generating process (or mechanism) likely becomes more complex than the observer’s tools or formal knowledge themselves Abrahão et al. (2023b). Only by adding a larger amount of (unpredictable) fresh data, i.e., new phenomena, not yet available to the observer in the first place would the observer hypothetically be capable of “bursting” this bubble. Only then, would the observer “realize” that the global optimum proposed by the model was not global at all.
@@ 2.1 No formal learning system can rule out demons in justified theories for lawlike universes
Let a hypothetical abstract entity, i.e., a ‘demon’, in a thought experiment correspond to a sufficiently large dataset (or to the underlying generative and governing agent for the set of every possible phenomena) which is an unpredictable deceiver that triggers the SBE for a given observer, i.e., a formal learning system. The above epistemological limitation is essentially a lower bound for the probability (given any arbitrary formal theory that defines the arbitrarily chosen probability measure) that such a demon exists. As the set of known and explainable phenomena becomes arbitrarily large, if Conditions I and II are satisfied, no one can rule out the possibility of those demons by endlessly decreasing the probability of their existence.
@@ 3 Interplay between Big Data, formal learning systems, and humanity
Leonelli proposes the notion of data journeys Leonelli (2014, 2023) to account for the material aspects, local idiosyncrasies, experimental settings, inferential processes, among other elements which influence the results in data-intensive research related to the processes of data achievement, integration, and use. Data journeys involve the process of de-contextualization of data from their local origin, such that it becomes digital and integrated in structured or semi-structured datasets that can be retrieved in other domains, allowing the possibilities of data reuse, reanalysis, recombination, repurposing, and repositioning Leonelli (2014); Collmann et al. (2016).
@@
An example of a data journey is the process of retrieval of information about amino acid sequences stored in genomic data infrastructures in order to design new forms of protein structures using freely available software for academic research Anishchenko et al. (2021). The storage of protein data in massive datasets, such as Uniprot, requires the process of standardization of the uploaded forms of proteins, being experimentally achieved or computationally designed. In addition, the dataset can impose some limitation in the retrieval of information. For example, a maximum of one hundred sequences per search. In the endeavor to cover as many species as possible, there could also be biases which favor the species that are well studied over fewer known organisms. Other kinds of biases may arise in the analysis phase itself as a limitation of statistics-based methods as is demonstrated in Uthamacumaran et al. (2024) when one tries to distinguish living from non-living matter.
@@
As presented in Cavassane et al. (2023), a potential problem illuminated by data journeys is the growing reliance on acknowledgeable datasets for decision-making processes in a software model which might be used to inform further research whose results might, in a feedback loop, serve as input to the data infrastructure. Our idea here is not to disregard the epistemic value of data integration, data modelling, and data retrieval, but to shed light on possible challenges of the growing automation of scientific practices by reflecting upon decision-making processes involved in data journeys. These challenges and a proposal on how to tackle them are discussed in Zenil et al. (2023b), in which new theoretical and technological tools are proposed in order to deal with the dynamics between AI-driven scientific discoveries and human scientists themselves once both will belong to the same knowledge production system in a closed-loop manner. We need to provide ways for making a given system accountable by its results in order to guide decision-making processes which can directly or indirectly affect society Alvarado (2023).
@@
In summary, data journeys constitute dynamic structures that, in association with computational tools, might guide scientific research and decision-making processes, promising new forms of scientific discovery. Once suitable for data journeys, data can grow in meaning by the process of recontextualization, with their original purpose being revised to fit a different research question or to redirect the course of investigation. Similar to the framework given by algorithmic information dynamics (AID) Zenil et al. (2020); Abrahão and Zenil (2022); Zenil et al. (2019a, 2018, b) which studies causal inference without the a priori separation between observers and objects, data journeys contrast the so-called ‘objective-oriented’ view of traditional scientific practice. One in which, by avoiding human interference and potential mistakes in scientific practice, we would eventually achieve neutral scientific results. This argument is succinctly summarized by Leonelli: “the paradox consists of the observation that, despite their epistemic value as ‘given’, data are clearly made” Leonelli (2014). Instead of this paradox suggesting an inexorable limit that science should acknowledge, novel methodologies should embed this paradigm shift in order to allow the investigation of both the objects and the observers themselves as interacting systems from the beginning and as part of the same theoretical and empirical framework Zenil et al. (2020); Abrahão and Zenil (2022); Zenil et al. (2023b).
@@
If we disregard, for the sake of the argument, the potential difficulties described above and imagine that a given dataset provides a reliable source for scientific evidence, the question to be reflected upon would be about the possibility to mechanically discriminate patterns which might unveil causal relationships among variables of interest from spurious correlations Floridi (2012); Smith (2020); Kitchin (2014); Calude and Longo (2017). A key criterion for the detection of causal relationships among variables is to discover the directionality of these variables, i.e., to uncover a causal chain (i.e., that rain causes the soil to be wet and not vice versa), or to show that two variables are the effect of the same cause (for instance, that thunder and lightning are caused by electric discharge). This reflects the Reichenbach principle: if two properties or kinds of events are probabilistically dependent, then they are causally connected in the sense that either one is a cause of the other (or vice versa), or both are effects of a common cause (where X is a cause of Y iff there leads a path of causal arrows from X to Y) Schurz (2008); Glymour et al. (1991).